Estimate to ±0.1 the pH value of the following solutions:

1. 0.043 M sodium hydrogen carbonate

2. 0.017 M sodium hydrogen tartrate

Blood pH is regulated to stay within the narrow range of 7.35 to 7.45, making it slightly basic. Which form of the following molecules is predominant at pH 7.4 (draw formula in ChemDraw or ChemSketch with appropriate protons shown):

1. Pyridine

2. Ethylamine

3. Acetic acid

4. Citric acid

5. Ibuprofen (pKa =4.91)

6. Glycine

7. Amphetamine (pKa =10.1)

8. Lactic acid
Answer #1

1) The reaction that occurs is:

HCO3- + H2O = H3O + + CO3-2

Where:

Ka = [H3O +] * [CO3-2] / [HCO3-] = X ^ 2 / 0.043 - X = 4.8x10 ^ -11

we clear and we have:

X ^ 2 + 4.8x10 ^ -11 X - 2.06x10 ^ -12 = 0

applying the equation of the second degree we have:

X = [H3O +] = 1.44x10 ^ -6 M

We calculate pH:

pH = -Log [H3O +] = - Log (1.44x10 ^ -6) = 5.84

2) The reaction that occurs is:

A- + H2O = HA + OH-

where:

Kb = [HA] * [OH-] / [A-] = X ^ 2 / 0.017 - X = 1.26x10 ^ -7

we cleared:

X ^ 2 + 1.26x10 ^ -7 X - 2.14x10 ^ -9 = 0

applying the equation of the second degree and we have:

X = [OH-] = 4.62x10 ^ -5 M

Calculating pOH and pH:

pOH = - Log (4.62x10 ^ -5) = 4.34

pH = 14 - 4.34 = 9.66

3) a) Pyridine

b) Ethylamine

c) Acetic acid

d) Citric acid

e) Ibuprofen

f) Glycine

g) Amphetamine

h) Lactic acid

If you like the answer, please help me with a positive rating, it would be great for me, regards.